From a collection of 10 different fonts, 3 different ones must be chosen to be used on a webpage: one for heading, one for standard text and one for quotations. Determine the number of different ways the fonts can be chosen and allocated, given that one of the ten fonts, MT Heading, is only appropriate for the heading.
The answer is 576.
Can someone please explain how to get this answer. Thank you.

Answer by greenestamps(13198)



If MT Heading is used for the heading, then, since the three fonts have to be different, there are 9 choices for the second font and 8 choices for the third; the total number of choices is 9*8=72.

If MT Heading is not used for the heading, then it can't be used for either text or quotations. So there are 9 choices for the header, then 8 choices for one of the others and 7 choices for the last; the total number of choices is 9*8*7=504. The total number of ways to choose the three fonts is 72+504=576.